OpenStudy (anonymous):

Put these in least to greatest -8/13 -3/4 -7/8 -2/9 -3/11

OpenStudy (anx):

The easiest method to do this is to turn the fractions into decimals. You have: -8/13 = -.615 -3/4 = -.750 -7/8= -.875 -2/9 = -.222 -3/11 = -.273 From least to greatest you have -2/9 -3/11 -8/13 -3/4 -7/8

OpenStudy (anonymous):

Thank You

OpenStudy (anonymous):

-2/9 , -3/11 , -8/13 , -3/4 , -7/8 @pinkglamx

OpenStudy (anonymous):

Need any more help @pinkglamx